# A Formal Comparison Between Chain-of-Thought and Latent Thought

### v1

@@ D.1.3 Self-Reducibility
On the event that all estimates in equation 104 are $(1\pm\varepsilon)$-accurate, we obtain
$\frac{1-\varepsilon}{1+\varepsilon}\,p(y_{i}\mid x,y_{<i})\;\leq\;\pi(y_{i}\mid x,y_{<i})\;\leq\;\frac{1+\varepsilon}{1-\varepsilon}\,p(y_{i}\mid x,y_{<i})\qquad\text{for all }y_{i}\in\Sigma.$ (106)
Choosing $\varepsilon\leq\tfrac{1}{2\alpha(|x|)+1}$ for any polynomial $\alpha(n)\in\mathsf{poly}(n)$ yields
$\frac{1+\varepsilon}{1-\varepsilon}\leq 1+\frac{1}{\alpha(|x|)},\qquad\frac{1-\varepsilon}{1+\varepsilon}\geq 1-\frac{1}{\alpha(|x|)},$
thus, we have an approximation to the true conditional distribution.
Finally, let $m(n)\in\mathsf{poly}(n)$ be such that $|y|\leq m(|x|)$ for all $y\in R(x)$. Applying a union bound over the $m(n)$ calls in Equation 104, if each call uses failure probability $\delta\leq\tfrac{\tfrac{1}{2}-\gamma}{m(|x|)}$ for any fixed $\gamma\in(0,\tfrac{1}{2})$, then the joint success event holds with probability at least $1-(|\Sigma|+1)\delta\geq\tfrac{1}{2}+\gamma$. Thus there exists an FPRAS ensuring
$\Pr\left[(1-\tfrac{1}{\alpha(|x|)})p(y_{i}\mid x,y_{<i})\;\leq\;\pi(y_{i}\mid x,y_{<i})\;\leq\;(1+\tfrac{1}{\alpha(|x|)})p(y_{i}\mid x,y_{<i})\right]\;\geq\;\tfrac{1}{2}+\gamma.$ (107)
By Lemma D.8, any PTM that realizes an FPRAS can be simulated by a log-precision CoT model with stochastic decoding, whose output distribution coincides exactly with that of the PTM. Hence, there exists a CoT model with polynomially many reasoning steps that achieves the $(\alpha,\gamma)$-weak probable approximation to the uniform distribution $U(R(x))$.
Proof of the Looped TF Part
Suppose, for contradiction, that there exists a Looped TF $\pi$ with polynomial loops, which, for every input $x$, achieves an $(\alpha,\gamma)$-weak probable approximation to the uniform distribution $U(R(x))$ with
$\alpha(|x|)\;\geq\;\frac{\max\{\log(1+\varepsilon),\;\log(1-\varepsilon)+\tfrac{1}{2}\}}{m(|x|)},$ (108)
where $m(n)\in\mathsf{poly}(n)$ satisfies $|y|\leq m(|x|)$ for any $y\in R(x)$, with $\varepsilon>0$ and constant $\gamma<1/2$.
By definition of the model, all randomness is injected only during decoding. The logits, and thus the conditional distributions $\pi(\cdot\mid x,y_{<i})$, are deterministically computable in time $\mathsf{poly}(|x|)$. Hence, the guarantee reduces to a deterministic bound: for every $i$ and prefix $y_{<i}$,
$(1-1/\alpha(|x|))\,p(y_{i}\mid x,y_{<i})\;\leq\;\pi(y_{i}\mid x,y_{<i})\;\leq\;(1+1/\alpha(|x|))\,p(y_{i}\mid x,y_{<i}).$ (109)
For a fixed string $\hat{y}\in R(x)$, define the polynomial-time computable quantity
$\widehat{Z}(x)\;\coloneq\;\Biggl(\prod_{i=1}^{|\hat{y}|}\pi(\hat{y}_{i}\mid x,\hat{y}_{<i})\Biggr)^{-1}.$ (110)
Since $p(y\mid x)=1/|R(x)|=\prod_{i}p(y_{i}\mid x,y_{<i})$, we have
$(1-1/\alpha(|x|))^{m}\;\leq\;\frac{\widehat{Z}(x)}{|R(x)|}\;\leq\;(1+1/\alpha(|x|))^{m}.$ (111)
By Lemma D.4, whenever $\varepsilon^{2}\leq m$,
$e^{\,\varepsilon-\tfrac{1}{2}}\;\leq\;\Bigl(1+\tfrac{\varepsilon}{m}\Bigr)^{m}\;\leq\;e^{\,\varepsilon}.$ (112)
Substituting $\varepsilon=\max\{\log(1+\varepsilon),\;\log(1-\varepsilon)+\tfrac{1}{2}\}$ into the inequality, we have
$(1-\varepsilon)\,|R(x)|\;\leq\;\widehat{Z}(x)\;\leq\;(1+\varepsilon)\,|R(x)|,$ (113)
i.e., a deterministic $(1\pm\varepsilon)$-approximation to $|R(x)|$ computable in time $\mathsf{poly}(|x|,1/\varepsilon)$. This constitutes an $\mathsf{FPTAS}$, contradicting the assumption that $\mathsf{FPTAS}\subsetneq\mathsf{FPRAS}$.
From the preceding argument, there exist a polynomial function $\alpha(n)\in\mathsf{poly}(n)$ and a constant $\gamma>0$ such that, for any input $x$, any Looped TF with polynomial loops has a first-step conditional $\pi(\cdot\mid x)$ that deviates from the target $p(\cdot\mid x)=U(R(x))$ on a set of symbols of non-negligible mass. Formally, define
$A\;\coloneq\;\Bigl\{\,a\in\Sigma\;\Bigm|\;\frac{\pi(a\mid x)}{p(a\mid x)}\leq 1-\frac{1}{\alpha(|x|)}\;\;\text{or}\;\;\frac{\pi(a\mid x)}{p(a\mid x)}\geq 1+\frac{1}{\alpha(|x|)}\Bigr\},$ (114)
and assume $p(A\mid x)\geq\gamma$. Let $q(y\mid x)\coloneq\prod_{k=1}^{m}\pi(y_{k}\mid x,y_{<k})$ be the model’s full distribution over strings. By data processing (marginalization cannot increase total variation),
$\|q(\cdot\mid x)-p(\cdot\mid x)\|_{\mathrm{TV}}\;\geq\;\|\pi(\cdot\mid x)-p(\cdot\mid x)\|_{\mathrm{TV}}\;=\;\frac{1}{2}\sum_{a\in\Sigma}\bigl|\pi(a\mid x)-p(a\mid x)\bigr|.$ (115)
For every $a\in A$ we have $\bigl|\pi(a\mid x)-p(a\mid x)\bigr|=p(a\mid x)\,\bigl|\tfrac{\pi(a\mid x)}{p(a\mid x)}-1\bigr|\geq\tfrac{1}{\alpha(|x|)}\,p(a\mid x)$. Hence
$\|q(\cdot\mid x)-p(\cdot\mid x)\|_{\mathrm{TV}}\;\geq\;\frac{1}{2}\sum_{a\in A}\frac{1}{\alpha(|x|)}\,p(a\mid x)\;\geq\;\frac{\gamma}{2\,\alpha(|x|)}.$ (116)
Therefore, the total variation distance from $U(R(x))$ is bounded below by $\gamma/(2\alpha(|x|))$ and thus cannot be made arbitrarily small. In particular, such a model cannot realize an FPAUS for $U(R(x))$. ∎
Appendix E Experimental Details
E.1 Parallelizable Tasks
E.1.1 Task Settings
@@
We define a sequence prediction task based on finite groups such as the symmetric group $S_{5}$. Given a sequence of group elements of length $k$, the model is required to output the cumulative products obtained by scanning the sequence from left to right. Formally, for an input sequence $(g_{1},g_{2},\ldots,g_{k})$, the target sequence is $(g_{1},\;g_{1}g_{2},\;g_{1}g_{2}g_{3},\;\ldots,\;g_{1}g_{2}\cdots g_{k}).$ We follow the setting of Merrill & Sabharwal (2025b).
@@
To ensure the dataset has balanced labels, we generate undirected graphs according to the Erdős–Rényi model (Erdos & Renyi, 1959) $G(n,p)$, where $n$ is the number of vertices and each possible edge is included independently with probability $p$. In the supercritical regime ($pn=c>1$), a single “giant” connected component emerges, occupying a fraction $s\in(0,1)$ of the vertices, which satisfies $s=1-e^{-cs}.$ Consequently, the probability that two uniformly random vertices are both in this component—and hence mutually reachable—is approximately $s^{2}$. To target a reachability probability of $1/2$, we set $s\approx\sqrt{\tfrac{1}{2}}\approx 0.707,\qquad c\approx\tfrac{-\ln(1-s)}{s}\approx 1.74,$ and thus $p=\tfrac{c}{n}\approx\tfrac{1.7}{n}.$ In practice, for each graph of size $n$ we fix $p=1.7/n$, which empirically yields $\Pr[\text{reachable}]\approx 50\%$ for $n\in[50,100]$, up to small fluctuations. Other random models (e.g., stochastic block models) can also be tuned to achieve such a balance, but the Erdős–Rényi choice offers the simplest parametrization. We follow the encoding scheme of Sanford et al. (2024a). The input to the model is serialized as a flat token sequence consisting of three parts:
$\underbrace{v_{0}\,v_{1}\,\cdots\,v_{n-1}}_{\text{vertices}}\ \underbrace{e_{1}\,e_{2}\,\cdots\,e_{m}}_{\text{edges}}\ \underbrace{s,t}_{\text{query}},$
where each vertex is denoted by a token $v_{i}$, each edge is represented as a pair “u,v” with $u<v$, and the final token “s,t” specifies the source–target pair for the reachability query.
@@
Following Feng et al. (2023), we generate expressions over integers modulo $r$ using the four operations ${+,-,\times,\div}$, where multiplication and division are defined via precomputed modular tables. To guarantee that each expression evaluates to a specific target value, we grow expressions backwards: starting from a sampled number, we iteratively replace it with a binary sub-expression that preserves its value under modular arithmetic. Different from Feng et al. (2023), we fix the modulus to $r=3$, as our focus lies in evaluating the reasoning over expressions rather than exploring the properties of each modular arithmetic system.
@@
The Edit Distance task requires computing the minimum number of edit operations needed to transform one string into another. The allowed operations are insertion, deletion, and replacement of a single character, and the objective is to predict the total edit distance given two input strings. To build the dataset, we follow Feng et al. (2023). We first generate two strings over a randomly sampled alphabet. The first string has a fixed length, while the second string is produced in two possible ways: with probability $0.4$, it is drawn as a random string of nearly the same length (within $\pm 3$ characters), and with probability $0.6$, it is derived from the first string by applying a given number of random edit operations. Each edit operation is chosen uniformly among deletion, replacement, and insertion. To avoid trivial cases, string pairs that are identical or whose lengths differ excessively are rejected and resampled.
E.1.2 Sequential Algorithms for CoT
@@
(1) Word problem: for this task, the CoT algorithm proceeds by sequentially scanning the token sequence and producing at each prefix the evaluation result of the expression step by step. Thus, the overall length of the CoT sequence matches the input length. (2) Graph connectivity: Following Bavandpour et al. (2025), the algorithm sequence is simply the trace of a breadth-first search (BFS) starting from the source $s$. At each step, the model emits the incident edges of the currently expanded node in the order they are visited. The sequence terminates as soon as the target $t$ is discovered. To implement this algorithm, we maintain a list (“scratchpad”) initialized with a dummy marker and the source, $(\texttt{N},s)$. We iterate through this list from left to right (i.e., queue order). Whenever the current node $u$ is expanded, we append to the end of the list all incident edges $(u,v)$ for neighbors $v$, optionally followed by a separator token $(u,\texttt{N})$. (3) Arithmetic expression evaluation: Following Feng et al. (2023), the CoT takes the fully expanded expression and repeatedly evaluates one innermost subexpression, writing down the simplified expression at each step until only a single numeral remains. For example, $2*(0+1)/2\to 2*1/2\to 2/2\to 1.$ The overall CoT sequence has quadratic length. (4) Edit distance: Following Feng et al. (2023), the CoT algorithm outputs the DP table entries in the same order they are computed, i.e., row by row from top-left to bottom-right (topological order). This yields a quadratic-length CoT sequence.
E.1.3 Training Configuration
@@
We trained all models using the AdamW optimizer with a linear learning rate schedule. The initial learning rate was set to $1\times 10^{-4}$ with a weight decay of $0.01$, and a batch size of $256$. Training was continued until the training loss plateaued. For Looped TFs, curriculum learning was applied to all tasks except edit distance: the input size was increased by $2$ for the word problem task, and by $4$ for the connectivity and arithmetic evaluation tasks. The model architecture was based on standard Transformers with an embedding dimension of $256$. We used $4$ attention heads, and varied the number of Transformer layers depending on the task: two layers for word problems, a single layer for connectivity, and time-modulated (Xu & Sato, 2025) model with a single layer for Looped TF on both arithmetic evaluation and the edit distance task. For CoT, we use the same configuration of the Transformer block.
E.2 Approximation Task
@@
To generate the dataset, we first construct a DNF formula $F$ by sampling $m$ clauses, each consisting of $w$ distinct literals over $n$ Boolean variables. Each literal is independently assigned to be either positive or negated. The formula is then serialized into a token sequence in which each clause is represented by its index together with variable–value pairs such as “$2=$ +1+1” or “$4=$ 1-1”. For Looped TFs, we prepare 100,000100{,}000 training samples and 1,0001{,}000 test samples. For CoT models, we instead generate an online dataset with the following structure. To train the CoT models, we simulate a single trial of the randomized counting algorithm of Karp & Luby (1983) as follows:
  1. 1.
    A clause $C_{1}$ is sampled with probability proportional to $2^{-|C_{j}|}$ (uniform in our experiments).
  2. 2.
    An assignment consistent with $C_{1}$ is generated by fixing the variables in the clause $C_{1}$ and randomly setting the remaining variables.
  3. 3.
    A clause $C_{2}$ is then selected uniformly at random, and the assignment is checked against it.
  4. 4.
    If the assignment satisfies $C_{2}$, the outcome token is set to Success; otherwise, Fail.
The sequence concatenates the serialized formula, the sampled clause, the full assignment, and the verification outcome, separated by <sep> tokens and terminated by <eos>. The CoT model is trained in an autoregressive manner, where prediction targets are defined by shifting the token sequence while masking out the formula description. We trained the models using the AdamW optimizer with a linear learning rate schedule. The initial learning rate was set to $1\times 10^{-4}$, with a weight decay of $0.01$. We used a batch size of $256$ (reduced to $32$ for the $1000$-loop setting) and trained for 10,00010{,}000 iterations.

### v2

@@ C.3 Self-Reducibility and Complexity of Approximate Counting
Conditioned on the event that all estimates in Equation 94 are $(1\pm\varepsilon)$-accurate, the multiplicative error is bounded by:
$\frac{1-\varepsilon}{1+\varepsilon}\leq\frac{\pi(y_{i}\mid x,y_{<i})}{p(y_{i}\mid x,y_{<i})}\leq\frac{1+\varepsilon}{1-\varepsilon}.$ (95)
To ensure the cumulative approximation error remains within $(1\pm\varepsilon^{\prime})$, we set the local precision to $\varepsilon\leq\frac{\varepsilon^{\prime}}{2+\varepsilon^{\prime}}$, which yields $\frac{1+\varepsilon}{1-\varepsilon}\leq 1+\varepsilon^{\prime}$ and $\frac{1-\varepsilon}{1+\varepsilon}\geq 1-\varepsilon^{\prime}$. To ensure the failure probability is at most $\delta^{\prime}$, we apply a union bound over the $m(n)$ generation steps and the $|\Sigma|$ calls per step. By setting the local confidence to $\delta\leq\frac{\delta^{\prime}}{m(n)(|\Sigma|+1)}$, the joint success event holds with probability at least $1-\delta^{\prime}$. Under this event, the CoT correctly simulates an FPRAS for $p(y_{i}\mid x,y_{<i})$ in total time $\text{poly}(n,1/\varepsilon^{\prime},\log(1/\delta^{\prime}))$. Finally, since CoT can represent an FPRAS by Lemma 4.3, it satisfies the requirements for the construction.
On the other hand, we show that latent thought cannot compute such an approximation. Suppose, for contradiction, that the model could compute the conditional distribution $\pi(y_{i}\mid x,y_{<i})$ to within a $(1\pm\varepsilon)$ relative error in a single step. We define the estimator for the total count $Z(x)=|R(x)|$ as:
$\widehat{Z}(x)\;\coloneq\;\Biggl(\prod_{i=1}^{m(n)}\pi(y_{i}\mid x,y_{<i})\Biggr)^{-1}.$ (96)
Since the true distribution satisfies $p(y\mid x)=1/|R(x)|=\prod_{i}p(y_{i}\mid x,y_{<i})$, the relative error of $\widehat{Z}(x)$ is governed by the product of local errors:
$(1+\varepsilon)^{-m(n)}|R(x)|\leq Z(x)\leq(1-\varepsilon)^{-m(n)}|R(x)|,$ (97)
By setting $\varepsilon\leq\frac{\varepsilon^{\prime}}{2m(n)}$, we apply Proposition C.14 and the properties of multiplicative error:
$(1+\varepsilon)^{-m(n)}\geq 1-m(n)\varepsilon\geq 1-\varepsilon^{\prime}/2>1-\varepsilon^{\prime},$ (98)
and for sufficiently small $\varepsilon$,
$(1-\varepsilon)^{-m(n)}\leq 1+2m(n)\varepsilon\leq 1+\varepsilon^{\prime}.$ (99)
Substituting these into Equation 96, we obtain:
$(1-\varepsilon^{\prime})|R(x)|\leq\widehat{Z}(x)\leq(1+\varepsilon^{\prime})|R(x)|.$ (100)
This implies that if the model could compute $\pi$ accurately, $\widehat{Z}(x)$ would constitute an FPTAS for $|R(x)|$. However, under the standard complexity-theoretic assumption that $\mathsf{FPTAS}\subsetneq\mathsf{FPRAS}$ for self-reducible relations, this yields a contradiction. ∎
Appendix D Experimental Details
D.1 Fundamental Algorithmic Reasoning Tasks
D.1.1 Task Settings
@@
We define a sequence prediction task based on finite groups such as the symmetric group $S_{5}$. Given a sequence of group elements of length $k$, the model is required to output the cumulative products obtained by scanning the sequence from left to right. Formally, for an input sequence $(g_{1},g_{2},\ldots,g_{k})$, the target sequence is $(g_{1},\;g_{1}g_{2},\;g_{1}g_{2}g_{3},\;\ldots,\;g_{1}g_{2}\cdots g_{k}).$ We follow the setting of (Merrill and Sabharwal, 2025b).
@@
To ensure that the reachability labels are approximately balanced, we generate undirected graphs according to the Erdős–Rényi model (Erdos and Renyi, 1959) $G(n,p)$, where $n$ is the number of vertices and each possible edge is included independently with probability $p$. In the supercritical regime ($pn=c>1$), a single “giant” connected component emerges, occupying a fraction $s\in(0,1)$ of the vertices, which satisfies $s=1-e^{-cs}.$ Consequently, the probability that two uniformly random vertices are both in this component—and hence mutually reachable—is approximately $s^{2}$. To target a reachability probability of $1/2$, we set $s\approx\sqrt{\tfrac{1}{2}}\approx 0.707,c\approx\tfrac{-\ln(1-s)}{s}\approx 1.74,$ and thus $p=\tfrac{c}{n}\approx\tfrac{1.7}{n}.$ In practice, for each graph of size $n$ we fix $p=1.7/n$, which empirically yields $\Pr[\text{reachable}]\approx 50\%$ for $n\in[50,100]$. We follow the encoding scheme of Sanford et al. (2024a). The input to the model is serialized as a flat token sequence consisting of three parts: $v_{0}\,v_{1}\,\cdots\,v_{n-1}\;e_{1}\,e_{2}\,\cdots\,e_{m}\;s,t$ where each vertex is denoted by a token $v_{i}$, each edge is represented as a pair “u,v” with $u<v$, and the final token “s,t” specifies the source–target pair for the reachability query.
@@
Following (Feng et al., 2023), we generate expressions over integers modulo $r$ using the four operations ${+,-,\times,\div}$, where multiplication and division are defined via precomputed modular tables. To guarantee that each expression evaluates to a specific target value, we grow expressions backwards: starting from a sampled number, we iteratively replace it with a binary sub-expression that preserves its value under modular arithmetic. Different from (Feng et al., 2023), we fix the modulus to $r=3$, as our focus lies in evaluating the reasoning over expressions rather than exploring the properties of each modular arithmetic system.
@@
The Edit Distance task requires computing the minimum number of edit operations needed to transform one string into another. The allowed operations are insertion, deletion, and replacement of a single character, and the objective is to predict the total edit distance given two input strings. To build the dataset, we follow (Feng et al., 2023). We first generate two strings over a randomly sampled alphabet. The first string has a fixed length, while the second string is produced in two possible ways: with probability $0.4$, it is drawn as a random string of nearly the same length (within $\pm 3$ characters), and with probability $0.6$, it is derived from the first string by applying a given number of random edit operations. Each edit operation is chosen uniformly among deletion, replacement, and insertion. To avoid trivial cases, string pairs that are identical or whose lengths differ excessively are rejected and resampled. Finally, the shorter string is always placed first to maintain a consistent input format. An example instance in the format is shown below: s v d h s s e e  …  v e | s h d s s s s  …  e s e <sep> 20 Here, the two input strings are separated by the token “|”, “<sep>” marks the end of the inputs, and the final number “20” denotes the computed edit distance.
D.1.2 Training Configuration
Configuration of chain of thought
@@ D.1.2 Training Configuration
For CoT models, training is performed with supervision of step-by-step algorithms. (1) Word problem: for this task, the CoT algorithm proceeds by sequentially scanning the token sequence and producing at each prefix the evaluation result of the expression step by step. Thus, the overall length of the CoT sequence matches the input length. (2) Graph connectivity: Following Bavandpour et al. (2025), the algorithm sequence is simply the trace of a breadth-first search (BFS) starting from the source $s$. At each step, the model emits the incident edges of the currently expanded node in the order they are visited. The sequence terminates as soon as the target $t$ is discovered. To implement this algorithm, we maintain a list (“scratchpad”) initialized with a dummy marker and the source, $(\texttt{N},s)$. We iterate through this list from left to right (i.e., queue order). Whenever the current node $u$ is expanded, we append to the end of the list all incident edges $(u,v)$ for neighbors $v$, followed by a separator token $(u,\texttt{N})$. (3) Arithmetic expression evaluation: Following (Feng et al., 2023), the CoT takes the fully expanded expression and repeatedly evaluates one innermost subexpression, writing down the simplified expression at each step until only a single numeral remains. For example, $2*(0+1)/2\to 2*1/2\to 2/2\to 1.$ The overall CoT sequence has quadratic length. (4) Edit distance: Following (Feng et al., 2023), the CoT algorithm outputs the DP table entries in the same order they are computed, i.e., row by row from top-left to bottom-right (topological order). This yields a quadratic number of steps in the input length.
Optimization and model details.
@@
We trained all models using the AdamW optimizer with a linear learning rate schedule. The initial learning rate was set to $1\times 10^{-4}$ with a weight decay of $0.01$, and a batch size of $256$. Training was continued until the training loss plateaued. For looped TFs, curriculum learning was applied to all tasks except edit distance: the input size was increased by $2$ for the word problem task, and by $4$ for the connectivity and arithmetic evaluation tasks. The model architecture was based on standard Transformers with an embedding dimension of $256$. We used $4$ attention heads, and varied the number of Transformer layers depending on the task: two layers for word problems, a single layer for connectivity, and time-modulated (Xu and Sato, 2025) model with a single layer for looped TF, to stabilize training, on both arithmetic evaluation and the edit distance task. For CoT, we use the same configuration of the Transformer block.
Uniform selection.
To estimate a lower bound on the number of reasoning steps required by CoT for each task, we first follow the procedure introduced in prior work (Bavandpour et al., 2025). Specifically, given a complete CoT trajectory consisting of $T$ intermediate reasoning steps, we construct shortened trajectories by uniformly selecting $k$ step indices from $\{1,\dots,T\}$, i.e., by taking a uniformly spaced subsequence of the original trajectory. Only the selected intermediate steps are used, while the remaining steps are removed. We then evaluate task performance as a function of $k$, as shown in Table 3.
Table 3: Results for CoT on parallelizable tasks trained with uniform selection.
Task $\bm{n}$ Chain of Thought
8 16 32 64
Word Problem 64 0.8 0.8 0.8 100.0
Graph Connectivity 32 81.0 81.4 83.6 100.0
Arithmetic Evaluation 16 41.0 41.5 41.2 82.5
Edit Distance 16 69.2 70.3 82.6 94.8
Stepwise internalization.
We adopt stepwise internalization proposed by (Deng et al., 2024), a curriculum-based training procedure that gradually removes CoT tokens and encourages the model to internalize intermediate reasoning within its hidden states. Starting from a model trained on full CoT trajectories, we progressively truncate intermediate reasoning tokens according to a predefined schedule and finetune the model at each stage. Specifically, when the CoT length is greater than $128$, we remove $16$ tokens per stage until the remaining CoT length reaches $128$. We then continue removing $8$ tokens per stage until the CoT length is reduced to $8$, training the model for $16$ epochs at each stage. The results for each fundamental algorithmic reasoning task are shown in Fig.  10.
Figure 10: Accuracy as a function of the CoT steps during stepwise internalization.
D.2 Approximate Counting and Approximate Sampling
D.2.1 Approximate Counting of DNF Formulas
@@
To generate the dataset, we first construct a DNF formula $F$ by sampling $m$ clauses, each consisting of $w$ distinct literals over $n$ Boolean variables. Each literal is independently assigned to be either positive or negated. The formula is then serialized into a token sequence in which each clause is represented by its index together with variable–value pairs such as “$2=$ +1+1” or “$4=$ 1-1”. For looped TFs, we prepare 100,000100{,}000 training samples and 1,0001{,}000 test samples. For CoT models, we instead generate an online dataset with the following structure. To train the CoT models, we simulate a single trial of the randomized counting algorithm of Karp and Luby (1983). The sequence concatenates the serialized formula, the sampled clause, the full assignment, and the verification outcome, separated by <sep> tokens and terminated by <eos>. CoT model is trained in an autoregressive manner, where prediction targets are defined by shifting the token sequence while masking out the formula description. We trained the models using the AdamW optimizer with a linear learning rate schedule. The initial learning rate was set to $1\times 10^{-4}$, with a weight decay of $0.01$. We used a batch size of $256$ (reduced to $32$ for the $1000$-loop setting) and trained for 10,00010{,}000 iterations. For inference, we count the total number of iterations used for summing output tokens (steps) in CoT across trials, and the number of loop iterations in the looped TF.
D.2.2 Approximate Sampling of Graph Colorings
We consider the problem of approximately sampling a proper $k$-coloring of a graph, also known as almost-uniform generation, a canonical randomized task closely related to $\mathsf{P}-$hard counting problems. Given an undirected graph $G=(V,E)$ with $n=|V|$ vertices and maximum degree $\Delta$, a proper $k$-coloring is an assignment of colors from $\{1,\dots,k\}$ to vertices such that no adjacent vertices share the same color. Let $\Omega_{k}(G)$ denote the set of all proper $k$-colorings of $G$. We restrict attention to graphs of bounded degree and assume $k\geq 2\Delta+1.$ Under this condition, classical results in approximate counting show that the number of proper $k$-colorings admits a FPAUS (Jerrum, 1995). The approximation relies on Markov Chain Monte Carlo (MCMC) sampling using Glauber dynamics for graph colorings. Starting from an arbitrary proper coloring, the Markov chain repeatedly selects a vertex uniformly at random and proposes to recolor it with a randomly chosen color, accepting the update only if the resulting coloring remains proper. When $k\geq 2\Delta+1$, this Markov chain is known to be rapidly mixing, converging to the uniform distribution over $\Omega_{k}(G)$ in polynomial time. For our experiments, we generate an undirected Erdős–Rényi random graph, where each edge is included independently with probability $p=\frac{1.7}{n},$ using a fixed random seed for reproducibility. For simplicity and ease of analysis, we focus on small graphs with $n=3$ and set the number of colors to $k=5$.
Each sample is generated by running $T$ steps of Glauber dynamics on the space $\Omega_{k}(G)$ of proper $k$-colorings. Starting from a greedy proper initialization, at each step we uniformly select a vertex and a color; the recoloring is accepted if and only if it preserves properness. The final state of the Markov chain is treated as an approximate sample from the uniform distribution over $\Omega_{k}(G)$. In addition to the final coloring, we optionally record the entire sequence of proposals and accept/reject outcomes for CoT and use it as supervision, which is not included for latent thought. To train sequence models, we serialize the graph structure, the initial coloring, and either the MCMC history or the final coloring into a single token sequence. We evaluate the distribution of solutions generated by the model rather than only solution accuracy. Given an input $x$, we draw $N$ independent samples from the model using ancestral decoding. Generation continues until an end-of-sequence (EOS) token is produced, and from each generated sequence we extract the final $n$ tokens immediately preceding the first EOS token, which encode a candidate solution. The resulting samples define an empirical distribution $\hat{P}$ over generated solutions via normalized occurrence counts. For each instance, the task provides an exact enumeration $\mathcal{Y}$ of all valid solutions. We define the reference distribution $P_{\mathrm{true}}$ as the uniform distribution over this solution set, assigning probability $1/|\mathcal{Y}|$ to each $y\in\mathcal{Y}$ and zero otherwise. We quantify the discrepancy between the empirical distribution $\hat{P}$ and the uniform reference distribution using the total variation distance $\frac{1}{2}\sum_{y\in\mathcal{Y}\cup\hat{\mathcal{Y}}}\left|\hat{P}(y)-P_{\mathrm{true}}(y)\right|.$ We trained the models using the AdamW optimizer with a linear learning rate schedule. The initial learning rate was set to $1\times 10^{-4}$, with a weight decay of $0.01$. We used a batch size of $256$ and trained for 5,0005{,}000 iterations. For inference, we measure the average number of steps (loops) per generation. We generate N=50, [^0] N=50{,}000 samples for CoT and N=10, [^0] N=10{,}000 samples for looped TFs. The resulting histograms of CoT outputs over the target support are shown in Figure 11.
Figure 11: Output distributions of CoT compared against the uniform target distribution.